# HugAgent: Evaluating LLMs in Simulating Human-Like Individual Reasoning on Open-Ended Tasks

### v1

@@ Appendix I Surveillance Camera Questionnaire
From human CBN corpora we estimate topic-specific sufficient statistics $\Theta_{t}=\{\mu_{n},\sigma_{n},\mu_{e},\sigma_{e},\alpha_{\text{imp}},\beta_{\text{imp}},\alpha_{\text{conf}},\beta_{\text{conf}},\mathcal{V}\}$: node/edge count moments $(\mu_{n},\sigma_{n}),(\mu_{e},\sigma_{e})$, Beta parameters for node importance and edge confidence, and a topic vocabulary $\mathcal{V}$ for label generation. A content hash over source JSON files ensures cache validity.
@@
$|V|\sim\text{TruncNorm}(\mu_{n},\sigma_{n},[n_{\min},n_{\max}])$, $|E|\sim\min\{\text{LogNormal}(\mu^{\prime}_{e},\sigma^{\prime}_{e}),{|V|(|V|-1)}/{2}\}$ with log-space moments chosen to match $(\mu_{e},\sigma_{e})$.
@@
We sample a demographic profile $d$ with correlated marginals: age bands, gender, education, income, housing status, employment, location, children, rent-burden. Simple rules induce a small stance prior $\delta(d)$ (e.g., renter $\Rightarrow+\delta$ on pro-development; older age $\Rightarrow-\delta$ on rapid change). We set $s_{0}\leftarrow\text{clip}(s_{0}+\delta(d),1,10)$.
@@ 8   Realize $(e_{j},b_{j-1},b_{j})$ to text; append to transcript;
This appendix provides a detailed user guide and representative use cases for Trace-Your-Thinking, our semi-structured chatbot system designed to elicit human reasoning at scale. We describe both participant-facing (user) and researcher-facing (admin) views, followed by system outputs and illustrative use cases. Our design emphasizes three goals: (i) lowering barriers for participants, (ii) giving researchers flexible and reliable control, and (iii) producing structured outputs that make reasoning analyzable at scale.
@@
Figure 13: Participant view while AI is processing questions and asks the participant to wait
@@
Participants can choose between typed responses and spoken input, enabling a think-aloud protocol that captures more spontaneous reasoning processes. Figure 10 illustrates the onboarding screen where both input modes are explained, while Figure 11 shows a participant actively using voice input to answer a question. Once the interview begins, participants can monitor their progress via a persistent progress bar (Figure 12), which reduces fatigue by making task completion transparent. During processing, the interface explicitly signals when the system is handling input (Figure 13), avoiding user confusion or dropout.
@@
Figure 19: Guiding question editor. Researchers can toggle tutorial vs. research questions, reorder them dynamically, and experiment with alternative protocols.
@@
$\text{DA}=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\Big[\,\operatorname{sgn}(\hat{y}_{i}-\hat{y}_{i}^{\text{prev}})=\operatorname{sgn}(y_{i}-y_{i}^{\text{prev}})\,\Big],$
where $\operatorname{sgn}(\cdot)$ returns +1+1 for increase, 1-1 for decrease, and 0 for no change.
Appendix R Proof-of-Concept Mitigation Demonstrations
To illustrate the feasibility of the principles outlined in §7, we conducted small-scale demonstrations. These are not intended as definitive solutions but as evidence of tractability.
Directional bias.
We applied salience-based pruning: only the top-$k$ most context-relevant spans (selected via cosine similarity with the intervention prompt) were retained. On GPT-4o, this reduced overshoot errors in surveillance by $\sim$9% (MAE from 1.83 to 1.66).
Domain sensitivity.
We tested simple offset correction by centering updates within each domain distribution. Qwen2.5-32B improved update accuracy in zoning from 58.4% to 61.2%.
Tail-driven failures.
Replacing MSE with a Huber-style loss in synthetic fine-tuning reduced extreme 7-point flips by $\sim$30%.
Takeaway.
These small interventions already yield measurable improvements, supporting our claim that HugAgent provides a diagnostic arena for principled mitigation. We invite future work to build upon these directions with more systematic methods.
Appendix S Use of LLM Usage

### v2

@@ Appendix I Surveillance Camera Questionnaire
From human CBN corpora we estimate topic-specific sufficient statistics $\Theta_{t}=\{\mu_{n},\sigma_{n},\mu_{e},\sigma_{e},\alpha_{\text{imp}},\beta_{\text{imp}},\alpha_{\text{conf}},\beta_{\text{conf}},\mathcal{V}\}$: node/edge count moments $(\mu_{n},\sigma_{n})$ and $(\mu_{e},\sigma_{e})$, Beta parameters for node importance and edge confidence, and a topic vocabulary $\mathcal{V}$ for label generation. A content hash over source JSON files ensures cache validity.
@@
We draw $|V|\sim\text{TruncNorm}(\mu_{n},\sigma_{n},[n_{\min},n_{\max}])$ and $|E|\sim\min\{\text{LogNormal}(\mu^{\prime}_{e},\sigma^{\prime}_{e}),{|V|(|V|-1)}/{2}\}$ with log-space moments chosen to match $(\mu_{e},\sigma_{e})$.
@@
We sample a demographic profile $d$ with correlated marginals: age bands, gender, education, income, housing status, employment, location, children, and rent-burden. Simple rules induce a small stance prior $\delta(d)$ (e.g., renter $\Rightarrow+\delta$ on pro-development; older age $\Rightarrow-\delta$ on rapid change). We set $s_{0}\leftarrow\text{clip}(s_{0}+\delta(d),1,10)$.
@@ 8   Realize $(e_{j},b_{j-1},b_{j})$ to text; append to transcript;
This appendix provides a detailed user guide and representative use cases for Trace-Your-Thinking, our semi-structured chatbot system designed to elicit human reasoning at scale. We describe both participant-facing (user) and researcher-facing (admin) views, followed by system outputs and illustrative use cases. We use open science practices as an example here. Our design emphasizes three goals: (i) lowering barriers for participants, (ii) giving researchers flexible and reliable control, and (iii) producing structured outputs that make reasoning analyzable at scale.
@@
Figure 13: Participant view while AI is processing questions, but the participant can still answer the following questions without waiting
@@
Participants can choose between typed responses and spoken input, enabling a think-aloud protocol that captures more spontaneous reasoning processes. Figure 10 illustrates the onboarding screen where both input modes are explained, while Figure 11 shows a participant actively using voice input to answer a question. Once the interview begins, participants can monitor their progress via a persistent progress bar (Figure 12), which reduces fatigue by making task completion transparent. During processing, the interface will show the processing status while still generating new questions and allow users to answer (Figure 13).
@@
Figure 19: Guiding question editor. Researchers can toggle tutorial vs. research questions, reorder them dynamically, and experiment with alternative protocols. They can also choose to skip some questions by changing the status.
@@
We define directional accuracy as a two-stage weighted metric: (1) detecting whether a belief change occurred, and (2) if so, correctly predicting the direction of change (increase, decrease, or no change). To better reflect the importance of directional reasoning, a higher weight is assigned to the second stage.
$\text{DirAcc}=\lambda\cdot\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}\!\left[\,\Delta\hat{y}_{i}=0\wedge\Delta y_{i}=0\,\right]+(1-\lambda)\cdot\frac{1}{|\mathcal{C}|}\sum_{i\in\mathcal{C}}\mathbf{1}\!\left[\,\operatorname{sgn}(\Delta\hat{y}_{i})=\operatorname{sgn}(\Delta y_{i})\,\right],$ (4)
where $\Delta\hat{y}_{i}=\hat{y}_{i}-\hat{y}_{i}^{\text{prev}}$ and $\Delta y_{i}=y_{i}-y_{i}^{\text{prev}}$ denote predicted and true belief changes, respectively. $\mathcal{C}=\{\,i\mid\Delta\hat{y}_{i}\neq 0\wedge\Delta y_{i}\neq 0\,\}$ is the set of samples where both predicted and true beliefs changed. We set $\lambda=0.3$ by default, placing greater emphasis on directional correctness.
This weighting reflects the intuition that correctly inferring the direction of belief change is more informative than merely detecting whether a change occurred. While the first stage (change detection) captures a coarse perceptual judgment that can often be guessed in noisy or stable settings, the second stage(directional correctness)reveals whether the model truly understands and reasons about belief dynamics. Hence, emphasizing the latter better reflects a model’s fidelity to human-like reasoning processes and its capacity to simulate belief evolution.
Unified Aggregate Score.
To provide a single comprehensive measure of model performance across both static and dynamic belief tasks, we define a unified score that integrates the belief state inference (BSI) and belief dynamic update (BDU) components into a normalized value within $[0,1]$.
Specifically, $S_{\text{BSI}}$ denotes the normalized accuracy score for static belief state inference, while $S_{\text{BDU}}$ aggregates multiple metrics from the belief dynamic update task, including tolerance accuracy, normalized MAE, and directional reasoning accuracy.
The unified score is computed as:
$\text{UnifiedScore}=\tfrac{1}{2}S_{\text{BSI}}+\tfrac{1}{2}\Big[\tfrac{1}{2}\big(\tfrac{1}{2}S_{\text{BDU-mae-norm}}+\tfrac{1}{2}S_{\text{BDU-acc}}\big)+\tfrac{1}{2}S_{\text{Directional-acc}}\Big].$ (5)
where each subscore $S_{\cdot}\in[0,1]$ represents a normalized evaluation metric. For MAE-based components, normalization is defined as:
$S_{\text{BDU-mae-norm}}=\max\!\big(0,\;\min(1,\;1-\tfrac{\text{MAE}}{\text{MAE}_{\max}})\big),$ (6)
where $\text{MAE}_{\max}$ denotes the task-specific upper bound of allowable error.
The unified score assigns equal weights to the state component ($S_{\text{BSI}}$) and the update component ($S_{\text{BDU}}$). Within the update branch, tolerance accuracy and MAE are equally weighted (0.5 each) to ensure a balanced consideration of robustness and precision. The directional reasoning component ($S_{\text{Directional-acc}}$) is assigned an equal weight (0.5) relative to the combined MAE and accuracy branch, reflecting its comparable importance in capturing belief-updating dynamics.
Q.2 Computation Details and Track Usage
Unless otherwise specified, all quantitative analyses and unified aggregate score (UAS) computations are performed on the human track, which serves as the primary evaluation benchmark due to its ecological validity and authentic reasoning diversity.
The synthetic track follows the same survey and interview protocol but is designed for auxiliary and extensibility testing. It provides a controlled setting for examining model sensitivity, scaling behavior, and cross-domain generalization under scripted causal belief networks (CBNs). While the human track grounds evaluation in real participant reasoning, the synthetic track extends the benchmark toward scalable stress testing and potential future applications in simulated social environments.
In practice, model predictions for both tasks—belief-state inference (BSI) and belief dynamic update (BDU)—are first computed independently. All metrics (accuracy, MAE, and directional accuracy) are averaged across the three domains (healthcare, surveillance, zoning) before aggregation into the unified score defined in Equation 5. Reported results in Section 4 and subsequent findings are therefore based on the human track unless explicitly noted otherwise.
Appendix R Use of LLM Usage